Galaxykawaii2:

Situation A: A hang glider descended 33.57 feet, and then ascended 14.98 feet. Situation B: A football team gained yards in one play, and then lost of a yard in the next play. Situation C: A hiker walked 3 1/6 miles yesterday and 2 1/2 miles today. Situation D: A worker earned $23.66 Monday and spent $15.84 Tuesday. Calculate the sum of each situation. Order the situations from least to greatest based on the numeric value of the sums.

Galaxykawaii2:

PLEASE HELP

logan215:

whats the question

jhonyy9:

for A. what mean descended and ascended ? use these

jhonyy9:

ohh who asked is offline now ...

Extrinix:

@Galaxykawaii2, can you give us the fractions for Situation B please.

Extrinix:

Situation A: Descending is going down, ascending is going up. Descends 33.57 feet from 0, \(0-33.57\) \(=-33.57ft\) Ascends 14.98 feet, NOT from 0, but from -33.57. \(-33.57+14.98\) \(=-18.59ft\) So it would be \(-18.59\) for Situation A.

Extrinix:

Situation B can not be answered due to: `no numbers/fractions`.

Extrinix:

Situation C: Hiked a certain amount of miles TOTAL. Hiked \(3\dfrac{1}{6}\) miles yesterday. Hiked \(2\dfrac{1}{2}\) miles today. So you would just add these together, BUT the fractions are different. How many times can 2 go into 6?, 3 times. So it would look like: \(3\dfrac{1}{6}\) and \(2\dfrac{3}{6}\) Now add them together: \(3+2+\dfrac{1}{6}+\dfrac{3}{6}\) \(=5\dfrac{4}{6}\) Now we need to reduce the fraction, it can be divided by 2. \(\dfrac{4}{6} \rightarrow \dfrac{2}{3}\) So, \(5\dfrac{2}{3}\) would be your answer for Situation C.

Extrinix:

Situation D: Earned and spent. Worker earned 23.66 on Monday. \(+23.66\) Worker spent 15.84 on Tuesday \(23.66-15.84\) \(=7.82\) So the worker still has $7.82 left.

Extrinix:

Now, we can't sort them because we don't have the numbers for B.